# Hypothesis Generation with Large Language Models

### v2

@@ 1 Introduction
Hypothesis generation drives scientific progress. Mendel’s hypothesis on allele pairs lays the foundation for modern genetics; Einstein’s hypothesis in general theory of relativity led to the prediction and subsequent confirmation of gravitational waves. In the context of language modeling, the hypothesis on scaling law inspires recent progress in large language models (LLMs) (Kaplan et al., 2020). Despite the importance of hypothesis generation, as Ludwig & Mullainathan (2024) point out, science has been curiously asymmetric. While many scientific publications present extensive formal and empirical evaluation of hypotheses, the generation of hypotheses happens off-stage by researchers. In order to generate novel hypotheses, researchers may read literature, analyze data, pick the brain of each other, and even “hallucinate” (see Kekulé’s discovery of the structure of the benzene molecule (Rothenberg, 1995)). Given the rise of large language models (Brown et al., 2020; Anthropic, 2023; OpenAI, 2023b), we examine their potential of providing much needed assistance in hypothesis generation in this work.
@@
To generate high-quality hypotheses with LLMs, we propose an algorithm inspired by the upper confidence bound algorithm in multi-armed bandits (Auer, 2002) (HypoGeniC, Hypothesis Generation in Context; see Figure 1). Given initial hypotheses generated from a small number of examples, we need to assess their quality and propose new hypotheses to address their deficiencies. To navigate this exploration-exploitation tradeoff, we introduce a reward function and evaluate the top $k$ hypotheses for each training example. We maintain a wrong example bank to capture the gap in knowledge of the hypotheses pool, and generate new hypotheses based on the wrong example bank to close the gap.
@@
It is important to emphasize that although the utility of hypotheses in assisting downstream classification serves as an indicator for LLMs’ ability to generate hypotheses, our goal is not to maximize the classification performance. Rather, our primary interest lies in the quality of the hypotheses. Thus, it is critical for the hypotheses to be interpretable beyond the LLM used to produce the hypotheses. We show that hypotheses generated by one LLM (e.g., GPT-3.5-turbo) can be used to make accurate inference by another LLM (e.g., Mixtral). On an out-of-distribution dataset for Deceptive reviews, we can even outperform the oracle fine-tuned RoBERTa. Such cross generalization provides strong evidence that we are able to generate hypotheses of high quality. Furthermore, through a qualitative analysis, our generated hypotheses not only confirm theories from existing literature but also provide new insights about the tasks. For instance, one novel hypothesis is that “reviews that mention personal experiences or special occasions, such as birthdays, anniversaries, or weddings, are more likely to be truthful”. We encourage future research on deception detection to explore these novel hypotheses.
@@
We have publicly released the code and data for HypoGeniC at https://github.com/ChicagoHAI/hypothesis-generation.
@@ 3.1 Tasks and Datasets
Deceptive review detection is an instance of deception detection, a widely studied phenomenon in psychology and other social sciences (Granhag & Vrij, 2005). This particular task (Deceptive reviews) requires distinguishing genuine reviews from fictitious ones (Ott et al., 2011), where human performance is about chance (Lai & Tan, 2019). The dataset includes 800 genuine reviews and 800 fictitious reviews for 20 hotels in Chicago.
@@ 4.3 Qualitative Analysis
For Deceptive reviews, we find that deceptive reviews are more likely to be emotional, use superlatives, or contain information that could not have been directly experienced. Similar findings are also found by previous studies on Deceptive reviews (Lai et al., 2020; Anderson & Simester, 2014; Ott et al., 2011; Li et al., 2014). For Tweet popularity, we discover that tweets that are concise, with specific or relevant hashtags, or with emotional tones are more likely to be retweeted more, aligning with prior studies (Tan et al., 2014; Gligorić et al., 2019). For we find that revealing something new or using vivid language and imagery can drive engagement from readers to click on headlines. Previous studies also find these rules apply to online news headlines (Banerjee & Urminsky, 2021; Sadoski et al., 2000).
@@
Using vivid language and imagery helps. Banerjee & Urminsky (2021)
@@
Given mixed evidence from previous literature on the effect of “reading ease” on headline clicks, Banerjee & Urminsky (2021) finds that reading ease negatively impacts click-through rates in hrough careful feature engineering. Consistent with this result, we found that the hypotheses that claim “straightforward” and “clear” writing to be indicative of higher click-through rates have relatively lower accuracies during training.
@@ 5 Additional Related Work
Increasing attention has been brought to the use of LLMs in social science research (Ziems et al., 2024; Kim & Lee, 2023, i.a.). Our experiments demonstrate the potential of LLMs in generating hypotheses for social science research to discover unknown knowledge in the data. Furthermore, our approach can be extended to natural sciences for general scientific discovery.
@@ Appendix D Qualitative Analysis on Generated Hypotheses
Deceptive reviews tend to be more exaggerated. Anderson & Simester (2014)
Truthful reviews tend to use more balanced and objective tone. Anderson & Simester (2014)
@@
Revealing something new helps. Banerjee & Urminsky (2021)
Using vivid language and imagery helps. Banerjee & Urminsky (2021)
Headlines with high intensity of emotions would be clicked more. Banerjee & Urminsky (2021)
Action-oriented headlines are clicked more. Banerjee & Urminsky (2021)

### v3

@@ Abstract
[…]
[1]
Hypothesis Generation with Large Language Models
Yangqiaoyu Zhou, Haokun Liu, Tejes Srivastava Hongyuan Mei & Chenhao Tan Department of Computer Science University of Chicago, Toyota Technological Institute at Chicago Chicago, IL 60637, USA {zhouy1,haokunliu,tejess,chenhao}@uchicago.edu,hongyuan@ttic.edu
@@ 1 Introduction
Hypothesis generation drives scientific progress. Mendel’s hypothesis on allele pairs lays the foundation for modern genetics; Einstein’s hypothesis in general theory of relativity led to the prediction and subsequent confirmation of gravitational waves. In the context of language modeling, the hypothesis on scaling law inspires recent progress in large language models (LLMs) (Kaplan et al., 2020). Despite the importance of hypothesis generation, as Ludwig and Mullainathan (2024) point out, science has been curiously asymmetric. While many scientific publications present extensive formal and empirical evaluation of hypotheses, the generation of hypotheses happens off-stage by researchers. In order to generate novel hypotheses, researchers may read literature, analyze data, pick the brain of each other, and even “hallucinate” (see Kekulé’s discovery of the structure of the benzene molecule (Rothenberg, 1995)). Given the rise of large language models (Brown et al., 2020; Anthropic, 2023; OpenAI, 2023b), we examine their potential of providing much needed assistance in hypothesis generation in this work.
@@
To generate high-quality hypotheses with LLMs, we propose an algorithm inspired by the upper confidence bound algorithm in multi-armed bandits (Auer, 2002) (HypoGeniC  [^1] We have publicly released the code and data for HypoGeniC at https://github.com/ChicagoHAI/hypothesis-generation., Hypothesis Generation in Context; see Figure 1). Given initial hypotheses generated from a small number of examples, we need to assess their quality and propose new hypotheses to address their deficiencies. To navigate this exploration-exploitation tradeoff, we introduce a reward function and evaluate the top $k$ hypotheses for each training example. We maintain a wrong example bank to capture the gap in knowledge of the hypotheses pool, and generate new hypotheses based on the wrong example bank to close the gap.
@@
It is important to emphasize that although the utility of hypotheses in assisting downstream classification serves as an indicator for LLMs’ ability to generate hypotheses, our goal is not to maximize the classification performance. Rather, our primary interest lies in the quality of the hypotheses. Thus, it is critical for the hypotheses to be interpretable beyond the LLM used to produce the hypotheses. We show that hypotheses generated by one LLM (e.g., GPT-3.5-turbo) can be used to make accurate inference by another LLM (e.g., Mixtral). On an out-of-distribution dataset for Deceptive reviews, we can even outperform the oracle fine-tuned RoBERTa. Such cross generalization provides strong evidence that we are able to generate hypotheses of high quality. Furthermore, through a qualitative analysis, our generated hypotheses not only confirm theories from existing literature but also provide new insights about the task. For instance, one novel hypothesis is that “reviews that mention personal experiences or special occasions, such as birthdays, anniversaries, or weddings, are more likely to be truthful”. We encourage future research on deception detection to explore these novel hypotheses.
@@ 3.1 Tasks and Datasets
Deceptive review detection is an instance of deception detection, a widely studied phenomenon in psychology and other social sciences (Granhag and Vrij, 2005). This particular task (Deceptive reviews) requires distinguishing genuine reviews from fictitious ones (Ott et al., 2011), where human performance is about chance (Lai and Tan, 2019). The dataset includes 800 genuine reviews and 800 fictitious reviews for 20 hotels in Chicago.
@@ 4.3 Qualitative Analysis
For Deceptive reviews, we find that deceptive reviews are more likely to be emotional, use superlatives, or contain information that could not have been directly experienced. Similar findings are also found by previous studies on Deceptive reviews (Lai et al., 2020; Anderson and Simester, 2014; Ott et al., 2011; Li et al., 2014). For Tweet popularity, we discover that tweets that are concise, with specific or relevant hashtags, or with emotional tones are more likely to be retweeted more, aligning with prior studies (Tan et al., 2014; Gligorić et al., 2019). For we find that revealing something new or using vivid language and imagery can drive engagement from readers to click on headlines. Previous studies also find these rules apply to online news headlines (Banerjee and Urminsky, 2021; Sadoski et al., 2000).
@@
Using vivid language and imagery helps. Banerjee and Urminsky (2021)
@@
Given mixed evidence from previous literature on the effect of “reading ease” on headline clicks, Banerjee and Urminsky (2021) finds that reading ease negatively impacts click-through rates in hrough careful feature engineering. Consistent with this result, we found that the hypotheses that claim “straightforward” and “clear” writing to be indicative of higher click-through rates have relatively lower accuracies during training.
@@ 5 Additional Related Work
Increasing attention has been brought to the use of LLMs in social science research (Ziems et al., 2024; Kim and Lee, 2023, i.a.). Our experiments demonstrate the potential of LLMs in generating hypotheses for social science research to discover unknown knowledge in the data. Furthermore, our approach can be extended to natural sciences for general scientific discovery.
@@ 7 Limitations
Acknowledgments
We thank the anonymous reviewers for their suggestions. We also thank members of the Chicago Human+AI Lab for their helpful comments. This work is supported in part by NSF grants IIS-2126602 and an award from the Sloan Foundation.
@@ Appendix D Qualitative Analysis on Generated Hypotheses
Deceptive reviews tend to be more exaggerated. Anderson and Simester (2014)
Truthful reviews tend to use more balanced and objective tone. Anderson and Simester (2014)
@@
Revealing something new helps. Banerjee and Urminsky (2021)
Using vivid language and imagery helps. Banerjee and Urminsky (2021)
Headlines with high intensity of emotions would be clicked more. Banerjee and Urminsky (2021)
Action-oriented headlines are clicked more. Banerjee and Urminsky (2021)